# Microlensing of non-singular black holes at finite size: a ray tracing approach

### v1

@@ VII.2 Picolensing
Figure 11: We sketch a typical picolensing setup: two observers record the luminosity of a distant source. If a gravitational lens is present in between the observers and the source, the perceived brightness will vary strongly with their relative distance $\delta_{\text{obs}}$, assuming the observers are close enough together and positioned in such a way that one observer
@@ VII.3 Further work
Black hole dark matter [54, 55] is constrained by lensing [56, 57, 58, 59, 60] and extragalactic gamma ray flux [61, 29], but to date the asteroid mass window remains open [62]. As new data is expected from Subaru/HSC and the Vera C. Rubin LSST, we plan to extend our ray tracing code to contribute towards further probing that open window in the context of non-singular black hole models.
@@ Appendix A Parametrized post-Newtonian framework at fourth order
$\mu_{\text{tot,3}}$ $=\frac{2}{3A_{1}^{4}\beta}\bigg[-6A_{2}^{3}+12A_{1}A_{2}A_{3}-6A_{1}^{2}A_{4}+A_{1}^{4}A_{2}(-2+3d^{2})\big]\,$, (100)
@@ H Hayward 11 0 0 $2\xi^{2}$ 11 11 11 $1-\tfrac{1}{4}\xi^{2}$
W Simpson–Visser 11 0 0 0 11 $\tfrac{1}{8}\xi^{2} 1+1$$\tfrac{1}{8}\xi^{2}$ $1+\tfrac{1}{8}\xi^{2}+\tfrac{3}{128}\xi^{4}$
Table 2: Parametrized post-Newtonian (PPN) expansion coefficients for the Schwarzschild metric, Hayward metric, Minkowski core metric, and Simpson–Visser metric, up to fourth order as defined in Eq. (24), using the notation $\xi=\ell/(GM)$. Note that for the Simpson–Visser metric a transformation to a new radius variable is required, cf. Eq. (22), where we find $A(\tilde{r})=1-2GM/\tilde{r}$ and $\tilde{B}=\tilde{r}^{2}/[(\tilde{r}-2GM)\sqrt{\tilde{r}^{2}-\ell^{2}}]$. Moreover, all of the above metrics feature $a_{1}=b_{1}=1$.

### v2

@@ VII.2 Picolensing
Figure 11: We sketch a typical picolensing setup: two observers record the luminosity of a distant source. If a gravitational lens is present in between the observers and the source, the perceived brightness will vary strongly with their relative distance $\delta_{\text{obs}}$, assuming the observers are positioned in such a way that one observer is within the Einstein ring radius and the other one outside of it.
@@ VII.3 Further work
Black hole dark matter [66, 67] is constrained by lensing [68, 69, 70, 71, 72] and extragalactic gamma ray flux [73, 36], but to date the asteroid mass window remains open [74]. As new data is expected from Subaru/HSC and the Vera C. Rubin Observatory, we plan to extend our ray tracing code to contribute towards further probing that open window in the context of non-singular black hole models.
@@ Appendix A Parametrized post-Newtonian framework at fourth order
$\mu_{\text{tot,3}}$ $=\frac{2}{3A_{1}^{4}\beta}\bigg[-6A_{2}^{3}+12A_{1}A_{2}A_{3}-6A_{1}^{2}A_{4}+A_{1}^{4}A_{2}(-2+3d^{2})\bigg]\,$, (113)
@@ H Hayward 11 0 0 $2\xi^{2}$ 11 11 11 $1-\tfrac{1}{4}\xi^{2}$
W Simpson–Visser 11 0 0 0 11 $\tfrac{1}{4}\xi^{2} 1+1$$\tfrac{1}{4}\xi^{2}$ $1+\tfrac{1}{4}\xi^{2}+\tfrac{1}{16}\xi^{4}$
Table 2: Parametrized post-Newtonian (PPN) expansion coefficients for the Schwarzschild metric, Hayward metric, Minkowski core metric, and Simpson–Visser metric, up to fourth order as defined in Eq. (24), using the notation $\xi=\ell/(GM)$. Note that for the Simpson–Visser metric a transformation to a new radius variable is required, cf. Eq. (22), where we find $A(\tilde{r})=1-2GM/\tilde{r}$ and $\tilde{B}=\tilde{r}^{3}/[(\tilde{r}-2GM)(\tilde{r}^{2}-\ell^{2})]$. Moreover, all of the above metrics feature $a_{1}=b_{1}=1$.